OpenStudy (anonymous):

can i get help please Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3.

OpenStudy (anonymous):

y-6=-3x-3 y-6=-3(x+1) y=-3x-3 3x+y=3

OpenStudy (anonymous):

the point-slope form of an equation is given by \[y-y _{1}=m(x-x _{1})\] where m is the slope and x1 and y1 represents the points that are known and lies on the line((-1,6) in this case) so the equation would be y-6=(x-(-1))(-3) y-6=(x+1)(-3) y-6=-3x-3